What is the slope of the line with the equation 2x + 3y + 6 = 0 ?
Found 2 solutions by Mathtut, checkley77:
Answer by Mathtut(3670)

write in slope intercept form y=mx+b where m is the slope and b is the y intercept
:
3y=-2x-6
:
y=(-2/3)x-2
:
m(slope)=-2/3

Answer by checkley77(12844)

2x + 3y + 6 = 0
3y=-2x-6
y=-2x/3-6/3
y=-2x/3-2 thus the slope is -2/3.